Consider the following data set 14.8 7.2 8.7 9.6 13.1 13.8 7.5 10.1 9.9 9.9 11.9 4.7 14.8 8.1 10.6 3.8 4.6 12.5 7.9 6.6 10.3 13.3 7.26.8 7.3 7.2 8.7 5.2 5.2 13.1 Click here for the Excel Data File a. Construct the frequency distribution using classes of 3 up to 5, 5 up to 7, etc. Class Frequency 3 up to 5 5 up to 7 7 up to 9 9 up to 11 11 up to 13 13 up to 15
b. Construct relative frequency, cumulative frequency, and cumulative relative frequency distributions. (Round Relative Frequency and Cumulative Relative Frequency to 2 decimal places.) 1.87 points Relative Cumulative Relative Frequency Cumulative Class 3 up to 5 5 up to 7 7 up to 9 9 up to 11 11 up to 13 13 up to 15 Frequency Freque eBook References c-1. How many of the observations are at least 7 but less than 9 Number of
c-2. How many of the observations are less than 9? Number of observations d-1. What percentage of the observations are at least 7 but less than 9? (Round your answer to the nearest whole percent.) centage of observations d-2. What percentage of the observations are less than 9? (Round your answer to the nearest whole percent.) of
Answer #1

-(a)
Frequency Distribution is as follows:

Class Frequency
3 up to 5 3
5 up to 7 4
7 up to 9 9
9 up to 11 6
11 up to 13 2
13 up to 15 6

(b)

Relative Frequency, Cumulative Frequency and Cumulative Relative Frequency distributions are as follows:

Class Relative Frequency Cumulative Frequency Cumulative Relative Frequency
3 up to 5 3/30 = 0.1000 3 3/30=0.1000
5 up to 7 4/30 = 0.1333 7 7/30=0.2333
7 up to 9 9/30 =0.3000 16 16/30=0.5333
9 up to 11 6/30=0.2000 22 22/30=0.7333
11 up to 13 2/30=0.0667 24 24/30=0.8000
13 up to 15 6/30=0.2000 30 30/30 = 1.0000

(c-1)

Number of observations at least 7 but less than 9 = 9

(c - 2)

Number of observations less than 9 = 3 + 4 + 9 = 16

(d - 1)

Percentage of observations at least 7 but less than 9 = (9/30)X100= 30 %

(d - 2)

Percentage of observations less than 9 = 3 + 4 + 9 = (16/30) X 100 = 53 %